Please help! The probability is 1 in 4,000,000 that a single auto trip in the US will result in a fatality. Over a lifetime, an average US driver takes 50,000 trips.
A) what is the probability of a fatal accident over a lifetime?
This is what was figured out:
1death/4,000,000 trips = x deaths /50,000
4,000,000x = 50,000
x=50,000/4,000,000
x=.0125 or 1.25%
My class is an online course, materials are printed. No hard book provided.

Answer by stanbon(48568)

Please help! The probability is 1 in 4,000,000 that a single auto trip in the US will result in a fatality. Over a lifetime, an average US driver takes 50,000 trips.
A) what is the probability of a fatal accident over a lifetime?
This is what was figured out:
1death/4,000,000 trips = x deaths /50,000
4,000,000x = 50,000
x=50,000/4,000,000
x=.0125 or 1.25%
------------------
Comment: Keep in mind that x is a probability, not a count.
x = (5000)*1/400,000 = 0.0125
=============================
Cheers,
Stan H.